a.) A doctor writes for amoxicillin 400mg/5 mL, give 2 tsp po bid for 10 days to treat an ear infection. The recommended dose for this indication is 45 mg/kg/dose. The child weights 39 lbs. What is an important question to ask the parents before beginning to process this antibiotic? Is the dose prescribed appropriate?

b.) Patient arrives to your Pharmacy with the following prescription: Oxycodone SR (Oxycontin) 80mg 1-2 tablets Q4h PRN pain. Do you have any inquiries for the pharmacist and/or provider regarding this prescription?

c.)An adult patient receives a prescription for Tylenol 500mg. The directions read T 2-4 tabs po q3-4h prn pain. Is the dose appropriate? Explain

d.) A prescription written for Novolog FlexPen U-100 has the following SIG: Inject 15 units subcutaneously before breakfast, 23 before lunch and 11 before dinner. What is the day supply for 2 boxes? (Each box has 5 X 3 mL pens)

e.) A prescription written for Timolol 0.5% 10mL 1gtt ou daily. What is the day supply of this prescription?
Answer #1

a) Before administering antibiotic, it is necessary to confirm with the parents whether the child is allergic to any medicines.

Step 1:

Recommended dose as per indication 45mg/Kg/dose.

Child weighs = 39lbs.

Child weight in Kg = 39/2.2 = 17.7

Child weight in Kg = 18 Kg(rounded off)

Step 2:

Recommended dose for the child = 45 X 18

= 810

The recommended dose for the Child is 810mg

Step 3:

The doctor ordered for 400mg/5mL 2 tsp.

5mL contains 400mg

therefore 2tsp(10mL) contains 800mg.

Hence, the ordered Dose is appropriate for the child.

b). Oxycodone SR is an opioid medication prescribed only when there is no alternative option. Because there is a high chance of drug abuse and misuse. The second thing Oxycodone is not recommended as PRN medicine. At last the dosage (80mg) is high and indicated to administer orally for every 12 hours. But the prescription shows 80mg 1-2 tab q4h. So the physician and pharmacist have to enquire about this.

C) No. this dose is not appropriate. Usually, Tylenol recommended a dose of 325mg 2 tablets can be administered every 4 to 6 hours interval in a day. The tablet should not exceed more than 10 in 24 hours. The maximum dose of Tylenol per day is 3000mg. But this prescription exceeds the maximum dose for a day. So it is not appropriate.

d)

Step 1:

Patient total units in a day = 15 +23 +11 = 49 units.

Step 2:

Total amount of units in each box

Novolog pre-filled pen 3mL contains 300Units

Each box contains 5 pens

The total amount of units in each box = 5 X 300 = 1500Units

Step 3:

Calculate the total amount of units in 2 boxes.

1 box contains 1500Units.

2 box contains 2 X 1500 = 3000Units.

Step 3:

Total day supply of two boxes = 3000/49

Total Day supply of two boxes = 61.22

The total daily supply of two boxes is for 61 days.

e)

Step 1:

Prescription says Timolol 1gtt ou daily.

That means instill 1 drop on each eye daily.

So the patient uses 2 drops per day.

Step 2:

The availability of Timolol is 10mL

1mL = 20 drops

10mL = 10 X 20 = 200drops.

Step 3:

Total supply days = 200 / 2

Total supply days = 100

The total supply day of this prescription is 100days.